# Bitangent surfaces and involutions of quartic surfaces

### v2

@@ 1. Introduction
Let $\mathbb{G}=G_{1}(\mathbb{P}^{3})$ be the Grassmannian of lines in $\mathbb{P}^{3}$. A surface $F\subset G_{1}(\mathbb{P}^{3})$ is called a congruence of lines. Its cohomology class $[F]$ in $H^{4}(\mathbb{G},\mathbb{Z})$ is determined by two numbers: the order $m$ and the class $n$. The order (resp. class) is the number of lines passing through a general point (resp. contained in a general plane) in $\mathbb{P}^{3}$. We will call the pair $(m,n)$ the bidegree of $F$. The number $m+n$ is equal to the degree of $F$ in the Plücker embedding $\mathbb{G}\hookrightarrow\mathbb{P}^{5}$ (cf. Dolgachev [3], Hudson [9], Jessop [11]).
@@
A surface $X$ in $\mathbb{P}^{3}$ defines a congruence of lines in $\mathbb{P}^{3}$ equal to the closure of the set of lines that touch $X$ at two distinct points. It is classically known as the bitangent surface, and we denote it by ${\mathrm{Bit}}(X)$. For a general $X$, its order and class are also classically known, and we will reproduce the calculation in this paper with the novelty of taking care of the case of characteristic $2$.
@@
For example, it is known that, for a general surface of degree $d$, the bidegree of the surface ${\mathrm{Bit}}(X)$ is equal to $(\frac{1}{2}d(d-1)(d-2)(d-3),\frac{1}{2}d(d-2)(d^{2}-9))$ [20, Volume 1, p. 281]. In particular, the bidegree of the bitangent surface of a general quartic surface is equal to $(12,28).$ Moreover, the surface is a smooth surface of degree $40$ in $\mathbb{P}^{5}$. It is of general type with $p_{g}=45$ and $K^{2}=360$ [24], [26].
@@
Although, for a general smooth quartic $X$, the surface ${\mathrm{Bit}}(X)$ is irreducible, for special quartic surfaces, even smooth ones, it can be reducible. It is known that, if a quartic surface $X$ is smooth and does not contain lines, then ${\mathrm{Bit}}(X)$ is smooth [1]. However, the surface could be reducible even when it is smooth but contains lines. For example, a general Cayley’s symmetroid quartic surface admits a smooth quartic model with reducible bitangent surface. One of its irreducible components is a Reye congruence of bidegree $(7,3)$ [2], [5, 7.4].
@@
Although admitting some mild singular points in $X$ doesn’t change the bidegree, it may drastically change the surface ${\mathrm{Bit}}(X)$, it even may become reducible. For example, when $X$ is realized as the focal surface of a congruence of lines of bidegree $(2,n)$ (for the focal surface, see [4, §2.1]), the surface ${\mathrm{Bit}}(X)$ is always singular although its singular points are ordinary double points. In the case $n=2$, the congruence of lines is a del Pezzo surface of degree $4$ and the focal surface is the famous $16$-nodal Kummer quartic surface $X$ (Kummer [15]), and the number of irreducible components of ${\mathrm{Bit}}(X)$ is equal 22, six of them are of bidegree $(2,2)$ and 16 of them are of bidegree $(0,1)$. They are planes of lines contained in one of the 16 planes intersecting $X$ along a conic, a trope plane.
@@
Table 1. Irreducible components of $\mathrm{Bit}(X)$ for Kummer quartic aurfaces $X$
Observe that $(12,28)$ is always a multiple of the sum of the bidegrees of irreducible components of the bitangent surface in characteristic 2. We do not know any explanation of this fact.
One reason for dropping of the order of ${\mathrm{Bit}}(X)$ in characteristic two is explained by the fact that the discriminant polynomial of a binary form is a square, and another is that irreducible components of bidegree $(1,1)$ appear instead of bidegree $(2,2)$. On the other hand, the class, being equal to the number of bitangent lines to a general plane section $H$ of $X$, also drops. It is equal to $7,4,2$, or $1$ if the Hasse-Witt invariant of $H$ is equal to $3,2,1$, or 00, respectively [19], [22]. We show that, in fact, the last case is not realized if $X$ is not a cone, i.e., a supersingular quartic curve can not be realized as a plane section of a non-conical quartic surface.
@@
The plan of the paper is as follows. In Section 2, we will reproduce Salmon’s proof for the formula of the bidegree of the bitangent surface of a general surface $X$ of degree $d$ in characteristic 00 (Theorem 2.1) and show it holds also for smooth surfaces or surfaces with rational double points (Corollary 2.3). In Section 3, we discuss the classical results about the bitangent surface of a Kummer quartic surface in characteristic $p=0$. They extend without change to the case $p\neq 2$. In Section 4, we discuss birational involutions of quartic surfaces. In Section 5, we give a proof (due to G. Kemper) of the fact that the discriminant of a binary form in the case $p=2$ is a square (Proposition 5.1). This reduces the order of congruences of bitangent lines to half. Finally, in the last three sections 6, 7 and 8 we determine the bidegrees of congruences of bitangent lines of Kummer quartics surfaces in characteristic two according to ordinary, $2$-rank 1 and supersingular, respectively (Theorems 6.5, 7.1, 8.1), and show that a supersingular quartic curve cannot be realized as a plane section of a non-conical quartic surface (Theorem 8.3).
@@
We would like to thank the referee for useful comments.
@@ 2. Generalities on the surface of bitangent lines
Let $X=V(F_{d})$ be a normal surface of degree $d$ and let ${\mathrm{Bit}}(X)$ be its bitangent surface, as it is defined in the introduction.
Theorem 2.1.
@@
$m\leq\frac{1}{2}d(d-1)(d-2)(d-3),\quad n=\frac{1}{2}d(d-2)(d^{2}-9).$
@@
We follow Salmon [20, Volume 2, p. 281]). Let $q$ be a general point in $\mathbb{P}^{3}$ and let $\ell=\langle q,q^{\prime}\rangle$ be the line containing $q$ and tangent to $X$ at some point $q^{\prime}=[x_{0},y_{0},z_{0},w_{0}]$. Without loss of generality, we may assume that $q=[0,0,0,1]$ and
@@
$f=\sum_{k+m=d}s^{k}t^{m}P_{q^{k}}(q^{\prime}),$
where $P_{v^{k}}(v^{\prime})=P_{v^{\prime m}}(v)$ is the value of $k$-th polar of $F$ with pole at $q$ at the vector $v^{\prime}$ (or, equivalently, the value of the totally polarized symmetric multilinear form defined by $f$ at $(v,\ldots,v,v^{\prime},\ldots,v^{\prime})$. Since $q^{\prime}\in X$, we get $P_{v^{0}}(v^{\prime})=F(v^{\prime})=0$. Moreover, because $\ell$ is tangent to $X$ at $q^{\prime}$, we obtain $P_{v}(v^{\prime})=0$. Thus, we can rewrite
@@
Recall that the discriminant polynomial $D(a_{0},\ldots,a_{d})$ of a binary form $\sum_{i=0}^{d}a_{i}u^{d-i}v^{i}$ of degree $d$ is a homogenous polynomial of degree $2(d-1)$. It is also a bi-homogeneous polynomial of bidegree $(d(d-1),d(d-1))$ with respect to the action of $\mathbb{G}_{m}^{2}$ via
$(a_{0},\ldots,a_{d})\mapsto(\lambda^{d}a_{0},\lambda^{d-1}\mu a_{1},\ldots,\lambda\mu^{d-1}a_{d-1},\mu^{d}a_{d}).$
@@
Thus, we obtain that the locus of points $q^{\prime}$ such that the line $\langle q,q^{\prime}\rangle$ is tangent to $X$ at two points including $q^{\prime}$ is contained in the intersection of hypersurfaces of degrees $d,d-1$, and $(d-2)(d-3)$. This implies that the expected number of bitangent lines passing through $q$ is equal to $\frac{1}{2}d(d-1)(d-2)(d-3)$.
@@
The class $n$ of ${\mathrm{Bit}}(X)$ is equal to the number of bitangent lines to a general plane section $H$ of $X$. Since $X$ is normal, it is a smooth plane curve of degree $d$. Their number is well-known classically, and it is equal to $\frac{1}{2}d(d-2)(d^{2}-9)$ (see [3, 5.5.1, the formula (5.33)]). ∎
@@
Let us see that the equality for $m$ is the expected order of ${\mathrm{Bit}}(X)$. This means that it holds for a general surface of degree $d$. To see this, we also introduce the congruence ${\mathrm{Flex}}(X)$, the closure of the locus of lines that intersect $X$ at some point with multiplicity $\geq 3$. Following the proof of the previous theorem, we find that the order of ${\mathrm{Flex}}(X)$ is equal to the degree of the complete intersection $X\cap P_{q}(X)\cap P_{q^{2}}(X)$. It is expected to be equal to $d(d-1)(d-2)$. The class of ${\mathrm{Flex}}(X)$ is equal to the number of flex tangents of $H$. It is equal to the number of intersection points of $H$ with its Hessian curve of degree $3(d-2)$. It is expected to be equal to $3d(d-2)$.
@@
Let ${\mathrm{pr}}_{q}:X\to\mathbb{P}^{2}$ be the projection of $X$ to a general plane. The ramification curve ${\mathrm{Rm}}(q)$ of the projection is equal to $X\cap P_{q}(X)$. Its degree is equal to $d(d-1)$ and its arithmetic genus is equal to $1+\frac{1}{2}d(d-1)(2d-5)$. The branch curve $B(q)$ of ${\mathrm{pr}}_{q}$ is equal to the projection of ${\mathrm{Rm}}(X)$, hence it is of degree $d(d-1)$. If ${\mathrm{Rm}}(q)$ is smooth, then it is isomorphic to the normalization of $B(q)$. A bitangent line containing the center of the projection is a secant line of ${\mathrm{Rm}}(X)$, hence its projection is a singular point of $B(q)$ with at least two different branches. For a quartic surface, it is an ordinary node. A flex line from ${\mathrm{Flex}}(X)$ containing $q$ is a tangent line to $P_{q}(X)$. Its projection is a singular point with at least two equal branches.
@@
Assume that $X$ is smooth or has only ordinary double points. Then, the bidegrees of ${\mathrm{Bit}}(X)$ and ${\mathrm{Flex}}(X)$ are equal to $(\frac{1}{2}d(d-1)(d-2)(d-3),\frac{1}{2}d(d-2)(d^{2}-9))$ and $(d(d-1)(d-2),3d(d-2))$, respectively
@@
We choose $q$ general enough such that no bitangent or flex line through $q$ passes through a singular point of $X$. The ramification curve ${\mathrm{Rm}}(q)$ of the projection ${\mathrm{pr}}_{q}$ is equal to the intersection $X\cap P_{q}(X)$ of $X$ with its polar with respect to $q$. We know that the jacobian ideal of an ordinary double point $x\in X$ is the maximal ideal $\mathfrak{m}_{X,x}$. This easily implies that $P_{q}(X)=\sum_{i=0}^{3}a_{i}\frac{\partial F_{d}}{\partial x_{i}}$ is smooth $x$ and, under the projection ${\mathrm{pr}}_{q}$, the double points of ${\mathrm{Rm}}(q)$ are mapped bijectively to ordinary double points of $B(q)$ different from the nodes and cusps coming from bitangent and flex lines.
Since ${\mathrm{Rm}}(q)$ is a complete intersection of degrees $(d,d-1)$, its arithmetic genus $p_{a}$ is equal to $p_{a}=1+\frac{1}{2}d(d-1)(2d-5)$. The projection ${\mathrm{pr}}_{q}$ defines a birational isomorphism between the curves ${\mathrm{Rm}}(q)$ and $B(q)$. The geometric genus $g$ of $B(q)$ is equal to $\frac{1}{2}(d(d-1)-1)(d(d-1)-2)-\delta-\kappa-\delta_{0}$, where $\delta$ is the number of bitangent lines, $\kappa$ is the number of flex lines, and $\delta_{0}$ is the number of nodes of ${\mathrm{Rm}}(q)$. This gives an inequality
@@
$\geq g=\frac{1}{2}(d(d-1)-1)(d(d-1)-2)-\delta-\kappa.$
Since $\delta\leq\frac{1}{2}d(d-1)(d-2)(d-3)$ and $\kappa\leq d(d-1)(d-2)$, we obtain the equality $\delta=\frac{1}{2}d(d-1)(d-2)(d-3)$ and $\kappa=d(d-1)(d-2)$.
One checks that this number coincides with the sum $\frac{1}{2}d(d-1)(d-2)(d-3)+d(d-1)(d-2)$ of the expected numbers of ordinary nodes and ordinary cusps of $B(q)$, which are the projections of the bitangent and flex line.
In this proof, we silently assumed that the curves ${\mathrm{Rm}}(q)$ and $B(q)$ are irreducible. One can treat the reducible case in a similar manner, by deviding the bitangents and flex lines into subsets corresponding to irreducible components of the curves. We leave it to the reader to finish the proof. ∎

### v3

@@ 1. Introduction
Let $G_{1}(\mathbb{P}^{3})$ be the Grassmannian of lines in $\mathbb{P}^{3}$. For a point in $G_{1}(\mathbb{P}^{3})$, the corresponding line in $\mathbb{P}^{3}$ is called a ray. A surface $F\subset G_{1}(\mathbb{P}^{3})$ is called a congruence of lines. Its algebraic class $[F]$ in the Chow ring $A^{\bullet}(G_{1}(\mathbb{P}^{3}))$ is determined by two numbers: the order $m$ and the class $n$. The order (resp. class) is the number of lines on $F$ passing through a general point (resp. contained in a general plane) in $\mathbb{P}^{3}$. We will call the pair $(m,n)$ the bidegree of $F$. The number $m+n$ is equal to the degree of $F$ in the Plücker embedding $G_{1}(\mathbb{P}^{3})\hookrightarrow\mathbb{P}^{5}$ (cf. [20, Chapter X, §2], [3], [11]).
@@
A surface $X$ in $\mathbb{P}^{3}$ defines a congruence of lines in $\mathbb{P}^{3}$ equal to the closure of the set of lines that are tangent to $X$ at two distinct points. It is classically known as the bitangent surface, and we denote it by ${\mathrm{Bit}}(X)$. For a general $X$, its order and class are also classically known, and we will reproduce the calculation in this paper with the novelty of taking care of the case of characteristic $2$.
@@
For example, it is known that, for a general surface of degree $d$, the bidegree of the surface ${\mathrm{Bit}}(X)$ is equal to $(\frac{1}{2}d(d-1)(d-2)(d-3),\frac{1}{2}d(d-2)(d^{2}-9))$ [19, Volume 1, p. 281]. In particular, the bidegree of the bitangent surface of a general quartic surface is equal to $(12,28).$ Moreover, the surface is a smooth surface of degree $40$ in $\mathbb{P}^{5}$. It is of general type with $p_{g}=45$ and $K^{2}=360$ [24], [27].
@@
Although, for a general smooth quartic $X$, the surface ${\mathrm{Bit}}(X)$ is irreducible, for special quartic surfaces, even smooth ones, it can be reducible. It is known that, if a quartic surface $X$ is smooth and does not contain lines, then ${\mathrm{Bit}}(X)$ is smooth [1]. However, the surface ${\mathrm{Bit}}(X)$ could be reducible even when $X$ is smooth but contains lines. For example, a general Cayley’s symmetroid quartic surface admits a smooth quartic model with reducible bitangent surface. One of its irreducible components is a Reye congruence of bidegree $(7,3)$ [2], [5, 7.4].
@@
Although admitting some mild singular points in $X$ doesn’t change the bidegree, it may drastically change the surface ${\mathrm{Bit}}(X)$, it even may become reducible. For example, when $X$ is realized as the focal surface of a congruence of lines of bidegree $(2,n)$ (the focal surface of a congruence $F$ of lines of order $n>1$ is a surface in $\mathbb{P}^{3}$ such that all rays of $F$ are its bitangents, see, for example, [20, Chapter X, §2], the surface ${\mathrm{Bit}}(X)$ is always singular although its singular points are ordinary double points. In the case $n=2$, the congruence of lines is a del Pezzo surface of degree $4$ and the focal surface is the famous $16$-nodal Kummer quartic surface $X$ [15], and the number of irreducible components of ${\mathrm{Bit}}(X)$ is equal to 22, six of them are of bidegree $(2,2)$ and 16 of them are of bidegree $(0,1)$. They are planes of lines contained in one of the 16 planes intersecting $X$ along a double conic, a trope plane.
@@
Table 1. Irreducible components of $\mathrm{Bit}(X)$ for Kummer quartic surfaces $X$
Observe that $(12,28)$ is a multiple of the sum of the bidegrees of irreducible components of the bitangent surfaces of ordinary Kummer surfaces in characteristic 2. We do not know any explanation of this fact.
One reason for the dropping of the order of ${\mathrm{Bit}}(X)$ in characteristic two is explained by the fact that the discriminant polynomial of a binary form is a square, and another is that irreducible components of bidegree $(1,1)$ appear instead of bidegree $(2,2)$. On the other hand, the class, being equal to the number of bitangent lines to a general plane section $H$ of $X$, also drops. It is equal to $7,4,2$, or $1$ if the Hasse–Witt invariant of $H$ is equal to $3,2,1$, or 00, respectively [18], [22].
@@
The plan of the paper is as follows. In Section 2, we will reproduce Salmon’s proof for the formula of the bidegree of the bitangent surface of a general surface $X$ of degree $d$ in characteristic 00 (Theorem 2.2) and show it holds also for smooth surfaces or surfaces with rational double points (Corollary 2.5). In Section 3, we discuss the classical results about the bitangent surface of a Kummer quartic surface in characteristic $p=0$. They extend without change to the case $p\neq 2$. In Section 4, we discuss birational involutions of quartic surfaces. In Section 5, we give a proof (due to G. Kemper) of the fact that the discriminant of a binary form in the case $p=2$ is a square (Proposition 5.3). This reduces the order of congruences of bitangent lines to half. Finally, in the last three sections 6, 7 and 8 we determine the bidegrees of congruences of bitangent lines of Kummer quartic surfaces in characteristic two according to ordinary, $2$-rank 1 and supersingular, respectively (Theorems 6.5, 7.1, 8.1).
@@
We would like to thank the referees for carefully reading the manuscript and for giving us many useful comments.
@@ 2. Generalities on the surface of bitangent lines
Let $X$ be a normal surface of degree $d\geq 4$. We keep this assumption during the whole paper. A line $\ell$ in $\mathbb{P}^{3}$ is called a bitangent line if either it is contained in $X$ or $X$ cuts out in $\ell$ a divisor $D=2a+2b+D^{\prime}$.
Let ${\mathrm{Bit}}(X)\subset G_{1}(\mathbb{P}^{3})$ be the variety of bitangent lines.
@@
The following fact is, of course, well known, but to be sure that it is true without assumption on the characteristic, we supply a proof.
Proposition 2.1.
An irreducible surface $S$ in $G_{1}(\mathbb{P}^{3})$ of bidegree $(m,0)$ (resp. $(0,n)$) is an $\alpha$-plane (resp. $\beta$-plane).
@@
Passing to the dual congruence of lines in the dual space of $\mathbb{P}^{3}$, it is enough to prove that any irreducible congruence of bidegree $(0,n)$ is an $\beta$-plane. Let $Z_{S}=\{(x,\ell)\in\mathbb{P}^{3}\times G_{1}(\mathbb{P}^{3}):x\in\ell\}$ be the restriction of the tautological projective bundle over $G_{1}(\mathbb{P}^{3})$ to $S$. Since $S$ is irreducible, $Z_{S}$ is an irreducible $3$-fold. Since $m=0$, its image $Y$ under the first projection is an irreducible subvariety of dimension $\leq 2$, the union of rays of $S$ that pass through some point in $Y$. Suppose $\dim Y=1$. Then, the fibers of $Z_{S}\to Y$ are $\alpha$-planes that sweep at least one plane. So, $Y$ cannot be a curve. Thus, $Y$ is a surface. The image of the fiber of $Z_{S}\to S$ over a point $y\in Y$ in $S$ is a curve contained in the $\alpha$-plane $\Omega(y)$ of lines through $y$. The corresponding rays sweep a cone with vertex at $y$ contained in $Y$. Thus, $Y$ is a cone at each of its points. Since $Y$ is reduced, this can happen only if $Y$ is a plane. So, all rays of $S$ are contained in a plane, and hence, $S$ is an $\beta$-plane. ∎
Proposition 2.2.
[…]
$1\leq m\leq\frac{1}{2}d(d-1)(d-2)(d-3),\quad 1\leq n=\frac{1}{2}d(d-2)(d^{2}-9).$
Proof.
We follow Salmon [19, Volume 2, p. 281]. Let $q$ be a general point in $\mathbb{P}^{3}$ and let $\ell=\langle q,q^{\prime}\rangle$ be the line containing $q$ and tangent to $X$ at some point $q^{\prime}=[x_{0},y_{0},z_{0},w_{0}]$. Without loss of generality, we may assume that $q=[0,0,0,1]$ and
@@
$f=\sum_{k+m=d}s^{k}t^{m}P_{v^{k}}(F)(v^{\prime}),$
where $P_{v^{k}}(F)(v^{\prime})$ is the value of the totally polarized symmetric multilinear form defined by $F$ at $(v,\ldots,v,v^{\prime},\ldots,v^{\prime})$. Geometrically, following the notation from [3, Chapter 1], the locus of zeros of $P_{v^{k}}(F)(v^{\prime})$ with fixed $v$ is the $k$-th polar $P_{q^{k}}(V(F))$ of the hypersurface $V(F)$. Recall from loc. cit. that the first polar hypersurface $P_{q}(V(F))$ is the locus of zeros of $\sum a_{i}\frac{\partial F}{\partial x_{i}}$, where $q=[a_{0},\ldots,a_{n}]$. Since $q^{\prime}\in X$, we get $P_{v^{0}}(F)(v^{\prime})=F(v^{\prime})=0$. Moreover, because $\ell$ is tangent to $X$ at $q^{\prime}$, we obtain $P_{v}(F)(v^{\prime})=0$. Thus, we can rewrite
@@
Recall that the discriminant polynomial $D(a_{0},\ldots,a_{n})$ of a binary form $\sum_{i=0}^{n}a_{i}s^{n-i}t^{i}$ of degree $n$ is a homogenous polynomial of degree $2(n-1)$. The polynomial $D(a_{0},\ldots,a_{n})$ is also a bi-homogeneous polynomial in variables $a_{0},\ldots,a_{n}$ of bidegree $(n(n-1),n(n-1))$ with respect to the action of $\mathbb{G}_{m}^{2}$ via
$(a_{0},\ldots,a_{n})\mapsto(\lambda^{n}a_{0},\lambda^{n-1}\mu a_{1},\ldots,\lambda\mu^{n-1}a_{n-1},\mu^{n}a_{n}).$
In other terms, it is a weighted homogeneous polynomial of degree $2(n-1)$ with the weights of the $a_{k}$’s equal to $k$.
@@
Applying this to the discriminant of the polynomial $g_{d-2}(s,t)$ we obtain that its discriminant $D(P_{v^{d}}(F)(v^{\prime}),\ldots,P_{v^{2}}(F)(v^{\prime}))$ is a polynomial of degree $\frac{1}{2}(d-2)(d-3)$ in $x_{0},y_{0},z_{0}$ and hence we obtain the locus of points $q^{\prime}$ such that the line $\langle q,q^{\prime}\rangle$ is tangent to $X$ at two points including $q^{\prime}$ is contained in the intersection of hypersurfaces of degrees $d,d-1$, and $(d-2)(d-3)$.
This implies that the expected number of bitangent lines passing through $q$ is equal to $\frac{1}{2}d(d-1)(d-2)(d-3)$.
@@
The class $n$ of ${\mathrm{Bit}}(X)$ is equal to the number of bitangent lines to a general plane section $H$ of $X$. Since $X$ is normal, the hyperplane section is a smooth plane curve of degree $d$. Their number is well-known classically, and it is equal to $\frac{1}{2}d(d-2)(d^{2}-9)$ (see [3, 5.5.1, formula (5.33)]).
[…]
To see that ${\mathrm{Bit}}(X)$ is a congruence of lines, i.e., each of its irreducible components is a surface, we consider the incidence variety
$M=\{(q,\ell)\in\mathbb{P}^{3}\times G_{1}(\mathbb{P}^{3}):\ell\in{\mathrm{Bit}}(X),q\in\ell\}.$
The fiber of the first projection consists of bitangent lines $\ell$ passing through $q$. By above, it consists of finitely many lines. This implies that each irreducible component of $M$ is of dimension $3$. Since the fibers of the second projections are lines, the image of each irreducible component of $M$ is a surface. ∎
We denote by ${\mathrm{Flex}}(X)\subset G_{1}(\mathbb{P}^{3})$ the variety of lines in $\mathbb{P}^{3}$ that either contained in $X$ or intersect $X$ at some point with multiplicity $\geq 3$. It intersects ${\mathrm{Bit}}(X)$ at the set of lines contained in $X$ and the set of bitangents that intersect $X$ at one point.
Proposition 2.3.
The expected bidegree of the congruence of lines ${\mathrm{Flex}}(X)$ is equal to $(m,n)$, where $m\leq d(d-1)(d-2)$ and $n\leq 3d(d-2)$.
Proof.
@@
Following the proof of the previous proposition, we find that the order $m$ of ${\mathrm{Flex}}(X)$ is equal to the degree of the reduced complete intersection $X\cap P_{q}(X)\cap P_{q^{2}}(X)$. We have $m\leq d(d-1)(d-2)$. The class $n$ of ${\mathrm{Flex}}(X)$ is equal to the number of flex tangents in a general plane section $H$ of $X$. It is equal to the number of intersection points of a plane section $H\cap X$ with its Hessian curve of degree $3(d-2)$. Clearly, $m\leq 3d(d-2)$. ∎
@@
Let ${\mathrm{pr}}_{q}:X\to\mathbb{P}^{2}$ be the projection of $X$ to a general plane. The ramification curve ${\mathrm{Rm}}(q)$ of the projection is equal to $X\cap P_{q}(X)$. Its degree is equal to $d(d-1)$. Using the well-known formula for the arithmetic genus of a complete intersection of hypersurfaces, we obtain that its arithmetic genus is equal to $1+\frac{1}{2}d(d-1)(2d-5)$. The branch curve $B(q)$ of ${\mathrm{pr}}_{q}$ is equal to the projection of ${\mathrm{Rm}}(X)$, hence it is of degree $d(d-1)$. If ${\mathrm{Rm}}(q)$ is smooth, then it is isomorphic to the normalization of $B(q)$. A bitangent line containing the center of the projection is a secant line of ${\mathrm{Rm}}(X)$, hence its projection has a singular point of $B(q)$ with at least two different branches. For a quartic surface, it is an ordinary node. A flex line from ${\mathrm{Flex}}(X)$ containing $q$ is a tangent line to $P_{q}(X)$. Its projection is a singular point with at least two equal branches.
@@
[…]
@@
We choose $q$ general enough such that no bitangent or flex line through $q$ passes through a singular point of $X$. The ramification curve ${\mathrm{Rm}}(q)$ of the projection ${\mathrm{pr}}_{q}$ is equal to the intersection $X\cap P_{q}(X)$ of $X$ with its polar with respect to $q$. We know that the jacobian ideal of an ordinary double point $x\in X$ is the maximal ideal $\mathfrak{m}_{X,x}$. This easily implies that $P_{q}(X)=V(\sum_{i=0}^{3}a_{i}\frac{\partial F_{d}}{\partial x_{i}})$ is smooth at $x$ and, under the projection ${\mathrm{pr}}_{q}$, the double points of ${\mathrm{Rm}}(q)$ are mapped bijectively to ordinary double points of $B(q)$ different from the nodes and cusps coming from bitangent and flex lines.
Since ${\mathrm{Rm}}(q)$ is a complete intersection of degrees $(d,d-1)$, its arithmetic genus $p_{a}$ is equal to $1+\frac{1}{2}d(d-1)(2d-5)$. The projection ${\mathrm{pr}}_{q}$ defines a birational isomorphism between the curves ${\mathrm{Rm}}(q)$ and $B(q)$. The geometric genus $g$ of $B(q)$ is equal to $\frac{1}{2}(d(d-1)-1)(d(d-1)-2)-\delta-\kappa-\delta_{0}$, where $\delta$ is the number of bitangent lines, $\kappa$ is the number of flex lines, and $\delta_{0}$ is the number of nodes of ${\mathrm{Rm}}(q)$. This gives an inequality
@@
$\geq g\geq\frac{1}{2}(d(d-1)-1)(d(d-1)-2)-\delta-\kappa$
that implies the inequality
(2.1) $\delta+\kappa\geq\frac{1}{2}d(d-1)(d-2)(d-3)+d(d-1)(d-2).$
Since $\delta$ is equal to the number of bitangents of $X$ dropped from $q$, we obtain $\delta$ is the order of ${\mathrm{Bit}}(X)$. Similarly, we get that $\kappa$ is the order of the ${\mathrm{Flex}}(X)$. Applying Propositions 2.2 and 2.3, we obtain $\delta\leq\frac{1}{2}d(d-1)(d-2)(d-3)$ and $\kappa\leq d(d-1)(d-2)$. It follows from (2.1) that $\delta=\frac{1}{2}d(d-1)(d-2)(d-3)$ and $\kappa=d(d-1)(d-2)$.
In this proof, we silently assumed that the curves ${\mathrm{Rm}}(q)$ and $B(q)$ are irreducible. One can treat the reducible case in a similar manner by dividing the bitangents and flex lines into subsets corresponding to irreducible components of the curves. We leave it to the reader to finish the proof. ∎